The stock of North American Dandruff Company is currently selling at $80 per share. The firm
pays a dividend of $2.50 per share.

a. What is the dividend yield?
b. If the firm has a payout rate of 50 percent, what is the firms P/E ratio?


Solution
Problem 18-10
Instructions

Enter formulas to calculate the requirements of this problem.

Information
Stock price$80
Dividend$2.50

a. What is the annual dividend yield?FORMULA in EXCEL

b. If the firm has a payout rate of 50 percent, what is the firms P/E ratio?FORMULA in Excel
Answer #1

a) dividend yield = 2.50/80 = 3.125%


b)earnins per share = 2.5/0.5 = 5


P/E ratio = 80/5 = 16
Answer #2
a. What is the annual dividend yield? FORMULA

Annual dividend yield = Current Dividend/Current Price = 2.50/80 = 3.125%

b. If the firm has a payout rate of 50 percent, what is the firm’s P/E ratio? FORMULA

P/E ratio = Price/Earnings = 5/80 = 6.25%

Earnings = Dividend/Payout ratio = 2.5/50% = 5